OpenStudy (anonymous):

(-13+15)-(3+2) divided by 6-12

OpenStudy (ddcamp):

Had a sign wrong: \[(-13+15)-(3+2) = 2 -5 = -3\] \[6-12 = -6\] \[\frac{ -3 }{ -6 } = \frac{ 1 }{ 2 }\]

OpenStudy (nincompoop):

you might want to resolve your numerator first resolve your denominator then apply division

OpenStudy (nincompoop):

why are you giving out answers, @DDCamp ? why????